OpenStudy (anonymous):

Visitors to a carnival can buy an unlimited-ride pass for $50 or an entrance-only pass for $20. In one day, 282 passes were sold for a total of $10,680. How many unlimited-ride passes were sold? 114 146 168 212 @texaschic101

OpenStudy (anonymous):

Its 168

OpenStudy (anonymous):

are u almost done kellie?

OpenStudy (texaschic101):

u = unlimited ride e = entrance only u + e = 282 --> u = 282 - e 50u + 20e = 10680 now we sub 282 - e in for u in the 2nd equation 50(282 - e) + 20e = 10680 14100 - 50e + 20e = 10680 -30e = 10680 - 14100 -30e = - 3420 e = 114 u + e = 282 u + 114 = 282 u = 282 - 114 u = 168 check.. 50u + 20e = 10680 50(168) + 20(114) = 10680 8400 + 2280 = 10680 10680 = 10680 (correct) unlimited tickets sold (u) = 168 entrance only tickets sold (e) = 114

OpenStudy (anonymous):

very good explanation :) thanks can u help me with some more

OpenStudy (anonymous):

are you ready for helping with some more do you have time to? @texaschic101

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

Simplify: 5 square root of 11 end root minus 12 square root of 11 end root minus 2 square root of 11 negative 33 square root of 9 negative 11 square root of 9 negative 9 square root of 11 negative 9 square root of 33

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

5 - 12 - 2 = 5 - 14 = - 9 the sq rt of 11 stays the same answer is -9 sq rt 11

OpenStudy (anonymous):

here is a snap shot this is weird

OpenStudy (anonymous):

thanks ready for another kellie?

OpenStudy (anonymous):

Which statement is true about the difference square root of 7 - square root of 28? It is rational and equal to -2. It is rational and equal to 0. It is irrational and equal to -2square root of 7. It is irrational and equal tosquare root of 7.

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

@DragonMaster .....not nice....how did you do that ? sorry LoneStarCuban....I am not good with square roots

OpenStudy (anonymous):

yeah thats not nice to do.. and can u tag someone that knows how ? @texaschic101

OpenStudy (anonymous):

@Preetha @Luigi0210

OpenStudy (anonymous):

Is the simplified form of 2square root of 3 • 2square root of 6 rational? Yes No @texaschic101

OpenStudy (anonymous):

@ganeshie8